# Deep Multimodal Learning with Missing Modality: A Survey

### v3

@@ 4.1.2. Distillation Based Methods
Representation-Based Distillation Methods transfer rich representations from the teacher model to help the student capture and reconstruct missing modality features. We classify them based on whether they use logits or intermediate features. Figure 11 illustrates this type of method. Response distillation methods focus on transferring teacher models’ logits to students, helping it mimic probability distributions. Wang et al. (Wang et al., 2020b) trained modality-specific teachers for missing modalities, then used their soft labels to guide a multimodal student. Hafner and Ban (Hafner and Ban, 2023) employed logits from a teacher model trained with optical data to supervise a reconstruction network for approximating missing optical features from radar data. Pramit et al. (Saha et al., 2024) proposed a Modality-Aware Distillation, leveraging both global and local teacher knowledge in a federated learning setting to help student models to learn how to handle missing modalities.
@@
Intermediate distillation methods align intermediate features between teacher and student models. Shen et al. (Shen and Gao, 2019) used Domain Adversarial Similarity Loss to align intermediate layers of the teacher and student, improving segmentation in missing modality settings. Zhang et al. (Zhang et al., 2023c) applied intermediate distillation in endometriosis diagnosis by distilling features from a TVUS-trained teacher to a student using MRI data.
@@
Process-Based Distillation Methods focus on overall distillation strategies, like Mean Teacher Distillation (MTD) (Tarvainen and Valpola, 2017) and self-distillation(Zhang et al., 2019). These methods emphasize procedural learning over direct representation transfer. MTD enhances stability by using the exponential moving average of the student model’s parameters as a teacher (Figure 12(a)). Chen et al. (Chen et al., 2024a) applied this to missing modality sentiment analysis, treating missing samples as augmented data. Li et al. (Li et al., 2022b) used MTD for Lidar-Radar segmentation, improving robustness against missing modalities.
@@
Self-distillation helps a model improve by learning from its own soft representations (Figure 12(b)). Wang et al. proposed the ShaSpec (Wang et al., 2023a), which utilizes distillation between modality-shared branches and modality-specific branches of available modalities. Based on ShaSpec, researchers proposed Meta-learned Cross-modal Knowledge Distillation (Wang et al., 2024b) to further weigh the importance of different available modalities to improve performance. Ramazanova et al. (Ramazanova et al., 2024) used mutual information and self-distillation for egocentric tasks, making predictions invariant to missing modalities. Shi et al. proposed PASSION (Shi et al., 2024), a self-distillation method designed to use the multi-modal branch to help other uni-modal branches of available modalities to improve multimodal medical segmentation performance with missing modality.
@@
Hybrid Distillation Methods combine various distillation approaches to improve student performance. For medical segmentation, Yang et al. (Yang et al., 2022) distilled teacher model knowledge (including logits and intermediate features) at every decoder layer, outperforming ACNet(Wang et al., 2021). Wang et al. (Wang et al., 2023d) introduced ProtoKD, which captures inter-class feature relationships for improved segmentation under missing modality conditions. Recently, CorrKD (Li et al., 2024a) leverages contrastive distillation and prototype learning to enhance performance in uncertain missing modality cases.
@@
Figure 13. Illustration of graph fusion methods. We set modality 2 (white circle in the graph) as the missing modalities. Features from modality 1 and 3 can be aggregated via this graph fusion method and fused features maintains the same dimension.
4.1.3. Graph Learning Based Methods
@@
Graph Fusion Methods integrate multimodal data using a graph structure (Figure 13), making them adaptable to various networks. For example, Angelou et al. (Angelou et al., 2019) proposed a method mapping each modality to a common space using graph techniques to preserve distances and internal structures. Chen et al. (Chen and Zhang, 2020) introduced HGMF, which builds complex relational networks using hyper edges that dynamically connect available modalities. Zhao et al. (Zhao et al., 2022) developed Modality-Adaptive Feature Interaction for Brain Tumor Segmentation, adjusting feature interactions across modalities based on binary existence codes. More recently, Yang et al. (Yang et al., 2023b) proposed a graph attention based Fusion Block to adaptively fuse multimodal features, using attention-based message passing to share information between modalities. These fusion methods can be flexibly inserted into any network for integrating multiple modalities.
@@
Figure 14. Illustration of graph neural network (GNN) based methods. We set modality 2 as the missing modalities. (a) Each modality first passes through DNN or builds its graph, and then they are uniformly processed using GNNs. (b) This method first complete the graph from missing-modality samples via approaches like FeatProp (Malitesta et al., 2024), then process it by GNNs.
Graph Neural Network Methods directly encode multimodal information into a graph structure, with GNNs used to learn and fuse this information. Early approaches (Wang et al., 2015) employed Laplacian graphs to connect complete and incomplete samples. Individual GNN methods (Figure 14(a)), such as DESAlign (Wang et al., 2024c), extract features using neural networks or GNNs and fuse them for prediction. Unified GNN methods ( Figure 14(b)) first complete the graph and then use GNNs for prediction, such as in Zhang et al.’s M3Care (Zhang et al., 2022), which uses adaptive weights to integrate information from similar patients. Lian et al. (Lian et al., 2023) proposed the Graph Completion Network, which reconstructs missing modalities by mapping features back to the input space. In recommendation systems, FeatProp(Malitesta et al., 2024) propagates known multimodal features to infer missing ones, while MUSE (Wu et al., 2024) represents patient-modality relationships as bipartite graphs and learns unified patient representations. In knowledge graphs, Chen et al. (Chen et al., 2023) introduced Entity-Level Modality Alignment to dynamically assign lower weights to missing/uncertain modalities, reducing the risk of misguidance in learning.
@@
4.1.4. Multimodal Large Language Models (LLM)
@@
Most MLLMs employ transformer-based modality encoders, such as CLIP, ImageBind (Girdhar et al., 2023), and LanguageBind (Zhu et al., 2023), which encode multimodal inputs into a unified representation space. Examples include BLIP-2 (Li et al., 2023b), which bridges the modality gap using a lightweight Querying Transformer, and LLaVA (Liu et al., 2024a), which enhances visual-language understanding with GPT-4-generated instruction data (Bubeck et al., 2023). These models are optimized for tasks like Visual Question Answering, Dialogue, and Captioning. Recent advancements extend output generation to multiple modalities, such as images. AnyGPT (Zhan et al., 2024) and NExT-GPT (Wu et al., 2023a) unify modalities like text, speech, and images using discrete representations and multimodal projection adaptors, enabling seamless multimodal interaction. CoDi (Tang et al., 2024) introduces Composable Multimodal Conditioning, allowing arbitrary modality generation through weighted feature summation. There are also some other methods that discard the modality encoders and use linear transformation directly, such as Fuyu (Bavishi et al., 2023) and OtterHD (Li et al., 2023c). Although MLLMs can flexibly handle any number of modalities, they have many disadvantages, such as the inconsistent multi-modal positional encoding, training difficulty, and high GPU resource requirements. Additionally, no specific MLLM benchmarks on missing modality problems have been proposed.
4.2. Model Combinations
@@
Figure 15. Two general architectures for ensemble methods. (a) The multimodal model ensemble methods contain $n$ three-modal DNNs and produce final predictions based on the outputs of all $n$ DNNs. (b) Each modality is processed by a unimodal DNN in unimodal model ensemble methods, where final predictions are produced by aggregating the outputs of all accessible unimodal DNNs.
4.2.1. Ensemble Methods
@@
Unimodal Model Ensemble Methods: The general architecture of this method is shown in Figure 15(b), where each modality is processed by a uni-modal model, and only the available modalities contribute to decision-making. In multimodal medical image diagnosis, early studies found uniformly weighted methods performed better than weighted averaging and voting approaches (Yuan et al., 2012). In multimodal object detection, Chen et al. (Chen et al., 2022) proposed a probabilistic ensemble method. This method does not require training and can flexibly handle missing modalities through probabilistic marginalization, demonstrating high efficiency in experiments. Li et al. (Li et al., 2023a) recently proposed Uni-Modal Ensemble with Missing Modality Adaptation, training models per modality and performing late-fusion training. Other approaches (Miech et al., 2018) calculate feature-based weights to fuse modalities, where weights reflect feature importance for final predictions.
@@
Hybird Methods: Dynamic Multimodal Fusion (DynMM) (Xue and Marculescu, 2023) uses gating mechanisms to select uni/multi-modal models dynamically. Recently, Cha et al. (Cha et al., 2024) proposed Proximity-based Modality Ensemble (PME) to use a cross-attention mechanism with an attention bias to integrate box predictions from different modalities. PME can adaptively combine box features from uni-/multi-model models and reduce noise in multi-modal decoding.
@@
Figure 16. General idea of dedicated training methods. Assuming that modality 2 is missing, dedicated methods select the sample-suited model from DNN models library for training and testing. In the figure, the three modalities can form a total of seven models with different modality combinations. For easy understanding, we use colored circles in purple rectangles to represent the modalities that the model can handle. In order to adapt to the missing modality 2 (light green circles), dedicated methods usually select the DNN model (purple rectangle) that can handle modality 1 (light orange circles) and modality 3 (light blue circles) for training and testing.
4.2.2. Dedicated Training Methods
@@
Dedicated training methods assign different tasks to specialized models. We show the general idea of those methods in Figure 16. KDNet (Hu et al., 2020) was the first dedicated method proposed to handle different combinations of missing modalities. It treats uni-modality specific models as student models, learning multimodal knowledge from the features and logits of a multimodal teacher model. Those trained uni-modals can be used for different missing modality cases. Following KDNet, ACNet (Wang et al., 2021) also utilizes this distillation method but introduces adversarial co-training, further improving the performance. Lee et al. (Lee et al., 2023b) proposed missing-modality-aware prompts to address missing modality problems based on the prompt learning by using input- and attention- level prompts for each kind of missing modality case. This method only needed 1% of the model parameters to be fine-tuned for downstream tasks. Similarly, some researchers (Reza et al., 2023) introduce different adapter layers for each missing modality case.
@@
Figure 17. General idea of discrete scheduler methods. The user uses text instructions to make LLM call the downstream task modules to complete the processing of data samples and feedback to the user. As long as there are sufficiently diverse downstream task modules, it is possible to handle a variety of tasks with different numbers of modalities.
4.2.3. Discrete Scheduler Methods
@@
In discrete scheduler methods (Figure 17), LLMs act as controllers, determining the execution order of different discrete steps broken down from the major task/instruction. While the LLM does not directly process multimodal data, it interprets language instructions and orchestrates task execution across uni- and multi-modal modules. This structured yet flexible approach is particularly effective for outputs requiring sequential tasks, enabling the system to handle any number of modalities and naturally addressing missing modality problems. For example, Visual ChatGPT (Wu et al., 2023b) integrates multiple foundation models with ChatGPT to enable interaction through text or images, allowing users to pose complex visual questions, provide editing instructions, and receive feedback within a multi-step collaboration framework. HuggingGPT (Shen et al., 2024) is an LLM-driven agent that manages and coordinates various AI models from Hugging Face. It leverages LLMs for task planning, model selection, and summarization to address complex multimodal tasks. ViperGPT (Surís et al., 2023) combines visual and language tasks into modular subprograms, generating and executing Python code for complex visual queries without additional training to achieve effective outputs. There are other similar discrete scheduler approaches, such as MM-REACT (Yang et al., 2023a) and LLaVA-Plus (Liu et al., 2023a).
@@
5. Methodology Discussion
Table 1. Comparison of deep multimodal learning with missing modality methods from four types and two aspects.
@@ 5. Methodology Discussion
Strategy Design Architecture Focused Attention-based Pros: Attention methods scale well; Distillation and graph learning are effective. Cons: Most attention methods demand high computational resources; Distillation and graph learning struggle with incomplete datasets and training complexity.
@@
In the above Section 3 and 4, we divide the existing MLMM methods into four types from the aspects of data processing and strategy design: modality imputation, representation-focused, architecture-focused, and model combinations. We also further subdivide these four types into twelve categories, exploring a fine-grained methodology taxonomy proposed by us. Table 1 summarizes the overall pros and cons of these methods. Generative and distillation methods are the most common approaches; they are easy to implement and deliver strong performance. With the rise of Transformers, Transformers methods have become more popular due to their larger receptive fields and parallelism. However, indirect-to-task generation methods and most distillation methods (except for some intermediate (Zhang et al., 2023c) and self-distillation methods (Wang et al., 2023a; Shi et al., 2024)) are currently unable to handle incomplete training datasets (means cannot access missing modality data during training). Below we provide a concise analysis of these twelve methods from two aspects and four types.
5.1. Data Processing Aspect

### v4

@@ 4.1.2 Distillation Based Methods
Representation-Based Distillation Methods transfer rich representations from the teacher model to help the student capture and reconstruct missing modality features. We classify them based on whether they use logits or intermediate features. figure 11 illustrates this type of method. Response distillation methods focus on transferring teacher models’ logits to students, helping it mimic probability distributions. Wang et al. (2020a) trained modality-specific teachers for missing modalities, then used their soft labels to guide a multimodal student. Hafner and Ban (2023) employed logits from a teacher model trained with optical data to supervise a reconstruction network for approximating missing optical features from radar data. A Modality-Aware Distillation was proposed by Saha et al. (2024), leveraging both global and local teacher knowledge in a federated learning setting to help student models to learn how to handle missing modalities.
@@
Intermediate distillation methods align intermediate features between teacher and student models. Shen and Gao (2019) used Domain Adversarial Similarity Loss to align intermediate layers of the teacher and student, improving segmentation in missing modality settings. Zhang et al. (2023b) applied intermediate distillation in endometriosis diagnosis by distilling features from a TVUS-trained teacher to a student using MRI data. Some researchers propose MultiPro (Xu et al., 2025) to learn the distribution of the missing modality by distilling unimodal knowledge into learnable unimodal prompts for each modality from all remaining samples available of this modality. Some also tackles the missing-modality problem in multimodal federated mobile sensing systems by introducing FedMobile (Liu et al., 2025a), a knowledge-contribution-aware federated learning framework that reconstructs missing features through cross-node multimodal knowledge transfer and robust aggregation.
@@
Process-Based Distillation Methods focus on overall distillation strategies, like Mean Teacher Distillation (MTD) (Tarvainen and Valpola, 2017) and self-distillation (Zhang et al., 2019). These methods emphasize procedural learning over direct representation transfer. MTD enhances stability by using the exponential moving average of the student model’s parameters as a teacher (figure 12(a)). Chen et al. (2024b) applied this to missing modality sentiment analysis, treating missing samples as augmented data. Li et al. (2022c) used MTD for Lidar-Radar segmentation, improving robustness against missing modalities.
@@
Self-distillation helps a model improve by learning from its own soft representations (figure 12(b)). Wang et al. proposed the ShaSpec (Wang et al., 2023a), which utilizes distillation between modality-shared branches and modality-specific branches of available modalities. Based on ShaSpec, researchers proposed Meta-learned Cross-modal Knowledge Distillation (Wang et al., 2024a) to further weigh the importance of different available modalities to improve performance. Ramazanova et al. (2024) used mutual information and self-distillation for egocentric tasks, making predictions invariant to missing modalities. Shi et al. proposed PASSION (Shi et al., 2024), a self-distillation method designed to use the multi-modal branch to help other uni-modal branches of available modalities to improve multimodal medical segmentation performance with missing modality.
@@
Hybrid Distillation Methods combine various distillation approaches to improve student performance. For medical segmentation, Yang et al. (2022) distilled teacher model knowledge (including logits and intermediate features) at every decoder layer, outperforming ACNet (Wang et al., 2021). Wang et al. (2023c) introduced ProtoKD, which captures inter-class feature relationships for improved segmentation under missing modality conditions. Recently, CorrKD (Li et al., 2024a) leverages contrastive distillation and prototype learning to enhance performance in uncertain missing modality cases.
@@
Figure 13: Illustration of graph fusion methods. We set modality 2 (white circle in the graph) as the missing modalities. Features from modality 1 and 3 can be aggregated via this graph fusion method and fused features maintains the same dimension.
4.1.3 Graph Learning Based Methods
@@
Graph Fusion Methods integrate multimodal data using a graph structure (figure 13), making them adaptable to various networks. For example, Angelou et al. (2019) proposed a method mapping each modality to a common space using graph techniques to preserve distances and internal structures. Chen and Zhang (2020) introduced HGMF, which builds complex relational networks using hyper edges that dynamically connect available modalities. Zhao et al. (2022) developed Modality-Adaptive Feature Interaction for Brain Tumor Segmentation, adjusting feature interactions across modalities based on binary existence codes. More recently, Yang et al. (2023a) proposed a graph attention based Fusion Block to adaptively fuse multimodal features, using attention-based message passing to share information between modalities. These fusion methods can be flexibly inserted into any network for integrating multiple modalities.
@@
Figure 14: Illustration of graph neural network (GNN) based methods. We set modality 2 as the missing modalities. (a) Each modality first passes through DNN or builds its graph, and then they are uniformly processed using GNNs. (b) This method first complete the graph from missing-modality samples via approaches like FeatProp (Malitesta et al., 2024), then process it by GNNs.
Graph Neural Network Methods directly encode multimodal information into a graph structure, with GNNs used to learn and fuse this information. Early approaches (Wang et al., 2015) employed Laplacian graphs to connect complete and incomplete samples. Individual GNN methods (figure 14(a)), such as DESAlign (Wang et al., 2024b), extract features using neural networks or GNNs and fuse them for prediction. Unified GNN methods ( figure 14(b)) first complete the graph and then use GNNs for prediction, such as in Zhang et al.’s M3Care (Zhang et al., 2022), which uses adaptive weights to integrate information from similar patients. Lian et al. (2023) proposed the Graph Completion Network, which reconstructs missing modalities by mapping features back to the input space. In recommendation systems, FeatPropMalitesta et al. (2024) propagates known multimodal features to infer missing ones, while MUSE (Wu et al., 2024b) represents patient-modality relationships as bipartite graphs and learns unified patient representations. In knowledge graphs, Chen et al. (2023b) introduced Entity-Level Modality Alignment to dynamically assign lower weights to missing/uncertain modalities, reducing the risk of misguidance in learning. Some work (Liu et al., 2020) addresses the missing-sensor (missing-modality) problem in topology-aware IoT systems by introducing a graph-based feature reconstruction module that uses neural message passing to recover features from available sensors. Some people Chen et al. (2023a) proposes a multi-graph convolutional framework for human activity recognition that explicitly handles asynchronous and incomplete IMU, magnetometer, and physiological signals, addressing missing-modality challenges in heterogeneous multi-sensor environments.
@@
4.1.4 Multimodal Large Language Models (LLM)
@@
Most MLLMs employ transformer-based modality encoders, such as CLIP, ImageBind (Girdhar et al., 2023), and LanguageBind (Zhu et al., 2023), which encode multimodal inputs into a unified representation space. Examples include BLIP-2 (Li et al., 2023b), which bridges the modality gap using a lightweight Querying Transformer, and LLaVA (Liu et al., 2024a), which enhances visual-language understanding with GPT-4-generated instruction data (Bubeck et al., 2023). These models are optimized for tasks like Visual Question Answering, Dialogue, and Captioning. Recent advancements extend output generation to multiple modalities, such as images. AnyGPT (Zhan et al., 2024) and NExT-GPT (Wu et al., 2023b) unify modalities like text, speech, and images using discrete representations and multimodal projection adaptors, enabling seamless multimodal interaction. CoDi (Tang et al., 2024) introduces Composable Multimodal Conditioning, allowing arbitrary modality generation through weighted feature summation. There are also some other methods that discard the modality encoders and use linear transformation directly, such as Fuyu (Bavishi et al., 2023) and OtterHD (Li et al., 2023a). Although MLLMs can flexibly handle any number of modalities, they have many disadvantages, such as the inconsistent multi-modal positional encoding, training difficulty, and high GPU resource requirements. Additionally, no specific MLLM benchmarks on missing modality problems have been proposed.
4.2 Model Combinations
@@
Figure 15: Two general architectures for ensemble methods. (a) The multimodal model ensemble methods contain $n$ three-modal DNNs and produce final predictions based on the outputs of all $n$ DNNs. (b) Each modality is processed by a unimodal DNN in unimodal model ensemble methods, where final predictions are produced by aggregating the outputs of all accessible unimodal DNNs.
4.2.1 Ensemble Methods
@@
Unimodal Model Ensemble Methods: The general architecture of this method is shown in figure 15(b), where each modality is processed by a uni-modal model, and only the available modalities contribute to decision-making. In multimodal medical image diagnosis, early studies found uniformly weighted methods performed better than weighted averaging and voting approaches (Yuan et al., 2012). In multimodal object detection, Chen et al. (2022) proposed a probabilistic ensemble method. This method does not require training and can flexibly handle missing modalities through probabilistic marginalization, demonstrating high efficiency in experiments. Li et al. (2023c) recently proposed Uni-Modal Ensemble with Missing Modality Adaptation, training models per modality and performing late-fusion training. Other approaches (Miech et al., 2018) calculate feature-based weights to fuse modalities, where weights reflect feature importance for final predictions.
@@
Hybird Methods: Dynamic Multimodal Fusion (DynMM) (Xue and Marculescu, 2023) uses gating mechanisms to select uni/multi-modal models dynamically. Recently, Cha et al. (2024) proposed Proximity-based Modality Ensemble (PME) to use a cross-attention mechanism with an attention bias to integrate box predictions from different modalities. PME can adaptively combine box features from uni-/multi-model models and reduce noise in multi-modal decoding.
@@
Figure 16: General idea of dedicated training methods. Assuming that modality 2 is missing, dedicated methods select the sample-suited model from DNN models library for training and testing. In the figure, the three modalities can form a total of seven models with different modality combinations. For easy understanding, we use colored circles in purple rectangles to represent the modalities that the model can handle. In order to adapt to the missing modality 2 (light green circles), dedicated methods usually select the DNN model (purple rectangle) that can handle modality 1 (light orange circles) and modality 3 (light blue circles) for training and testing.
4.2.2 Dedicated Training Methods
@@
Dedicated training methods assign different tasks to specialized models. We show the general idea of those methods in figure 16. KDNet (Hu et al., 2020) was the first dedicated method proposed to handle different combinations of missing modalities. It treats uni-modality specific models as student models, learning multimodal knowledge from the features and logits of a multimodal teacher model. Those trained uni-modals can be used for different missing modality cases. Following KDNet, ACNet (Wang et al., 2021) also utilizes this distillation method but introduces adversarial co-training, further improving the performance. Lee et al. (2023b) proposed missing-modality-aware prompts to address missing modality problems based on the prompt learning by using input- and attention- level prompts for each kind of missing modality case. This method only needed 1% of the model parameters to be fine-tuned for downstream tasks. Similarly, some researchers (Reza et al., 2023) introduce different adapter layers for each missing modality case.
@@
Figure 17: General procedure of discrete scheduler methods. The user types text instructions to make an LLM call and schedule the selected downstream task modules to process data samples and return results to the user. As long as there are sufficiently diverse downstream task modules, it is possible to handle the same downstream tasks with different numbers of modalities.
4.2.3 Discrete Scheduler Methods
@@
In discrete scheduler methods (figure 17), LLMs act as controllers, determining the execution order of different discrete steps broken down from the major task/instruction. While the LLM does not directly process multimodal data, it interprets language instructions and orchestrates task execution across uni- and multi-modal modules. This structured yet flexible approach is particularly effective for outputs requiring sequential tasks, enabling the system to handle any number of modalities and naturally addressing missing modality problems. For example, Visual ChatGPT (Wu et al., 2023a) integrates multiple foundation models with ChatGPT to enable interaction through text or images, allowing users to pose complex visual questions, provide editing instructions, and receive feedback within a multi-step collaboration framework. HuggingGPT (Shen et al., 2024) is an LLM-driven agent that manages and coordinates various AI models from Hugging Face. It leverages LLMs for task planning, model selection, and summarization to address complex multimodal tasks. ViperGPT (Surís et al., 2023) combines visual and language tasks into modular subprograms, generating and executing Python code for complex visual queries without additional training to achieve effective outputs. There are other similar discrete scheduler approaches, such as MM-REACT (Yang et al., 2023b) and LLaVA-Plus (Liu et al., 2023c).
@@
5 Methodology Discussion
Table 1: Comparison of deep multimodal learning with missing modality methods from four types and two aspects.
@@ 5 Methodology Discussion
Strategy Design Architecture Focused Attention-based Pros: Attention methods scale well; Distillation and graph learning are effective. Cons: Attention methods often demand large computation & datasets; Distillation and graph learning struggle with incomplete datasets and training complexity.
@@
In the above section 3 and section 4, we divide the existing MLMM methods into four types from the aspects of data processing and strategy design: modality imputation, representation-focused, architecture-focused, and model combinations. We also further subdivide these four types into twelve categories, exploring a fine-grained methodology taxonomy proposed by us. Table 1 summarizes the overall pros and cons of these methods. Generative and distillation methods are the most common approaches; they are easy to implement and deliver strong performance. With the rise of Transformers, Transformers methods have become more popular due to their larger receptive fields and parallelism. However, indirect-to-task generation methods and most distillation methods (except for some intermediate (Zhang et al., 2023b) and self-distillation methods (Wang et al., 2023a; Shi et al., 2024)) are currently unable to handle incomplete training datasets (means cannot access missing modality data during training). Below we provide a concise analysis of these twelve methods from two aspects and four types.
5.1 Data Processing Aspect